# PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

### v1

@@ 3.3 Tower Reuse
In CKKS, the Residue Number System (RNS) moduli are typically chosen as primes close to the scaling factor $\Delta$ (see §A for an overview of LFHE). This choice is motivated by the fact that, after each homomorphic multiplication, the scaling factors of the operands are multiplied, producing a ciphertext with scale $\Delta^{2}$. To control this growth, the rescale operation reduces the scale back to $\Delta$ by removing one modulus from the modulus chain. However, rescaling introduces approximation errors: the further the dropped modulus deviates from $\Delta$, the larger the error. These errors may accumulate and amplify through subsequent homomorphic operations. A larger $\Delta$ is desirable for two reasons: (i) it provides higher precision for CKKS encoding and (ii) the RNS moduli are relatively closer to the scaling factor, reducing approximation errors.
@@
The drawback is that a larger $\Delta$ requires larger RNS moduli, which in turn increases the ciphertext modulus $Q=\prod_{i=0}^{\mathcal{L}}q_{i}$, where $\mathcal{L}$ is the number of levels and $q_{i}$ are the individual RNS primes. Since the security level depends on the ratio $N/Q$ (with $N$ the polynomial degree), a larger $Q$ reduces security. One way to increase security is to increase $N$, but this slows computations due to larger polynomials. Thus, for a fixed security level and polynomial degree, there is a maximum allowable $Q$. Consequently, for an application-defined multiplicative depth $\delta$, this imposes a bound on the maximum scaling factor $\Delta$. In practice, when $\delta$ is large, as is common in deep learning models, the resulting $\Delta$ is small. This has two adverse effects, opposite to the benefits of a large $\Delta$: lower numerical precision and larger approximation errors, as moduli $q_{i}$ are relatively farther from $\Delta$.
@@
Proposed Solution. We introduce a more general method for determining the RNS moduli by introducing sublevels within each level. Instead of enforcing $q_{i}\approx\Delta$, we allow $q_{i}\approx\Delta^{\ell}$, where $\ell$ denotes the number of sublevels. Rescaling is performed only when a ciphertext scale exceeds the sublevel capacity of the modulus to be dropped. Formally, $\lambda(x)>\lambda(q_{i})\implies\downarrow x$, where $\downarrow$ denotes the rescale operation, and $\lambda(x)=\left\lfloor\log_{\Delta}{\Delta_{x}}\right\rceil$ returns the sublevel of a ciphertext, plaintext, or modulus: with $\Delta_{x}$ referring to the scale of $x$ and $\Delta$ being the default scale defined by the encryption parameters. See §E.3 for an example of this technique.
This method allows smaller $\Delta$ values while maintaining larger $q_{i}$, mitigating rescaling errors. Moreover, fewer effective levels (moduli) reduce error amplification, and since homomorphic operations are applied across all moduli, reducing their number improves computational efficiency.
@@ 3.4 Impact on RNS Levels
The combination of our polynomial activation function with the structural optimizations proposed in this section reduces the number of RNS levels $\mathcal{L}$ required for multiplication compared to PILLAR (Diaa et al., 2024), which previously achieved the lowest $\mathcal{L}$: from 87 to 18 for ResNet-18, from 97 to 20 for ResNet-20, and from 157 to 32 for ResNet-32. A complete analysis of the contribution of each technique to the reduction in required levels for ResNet models is provided in §E.4.
4 Codesign Techniques
@@ 4.1 Data Layout
Data layout defines how model inputs and weights are mapped into FHE ciphertext and plaintext slots during encoding. In this work, we adopt the HW layout (Dathathri et al., 2019), which assigns
Figure 1: Illustration of HW layout for convolution on a $3\times 3\times 3$ input and two $3\times 2\times 2$ filters, configured with padding 0 and stride 1.
each input channel to a separate ciphertext and fills its slots with spatial values indexed by two–dimensional coordinates. Unlike the naive layout (Gilad-Bachrach et al., 2016), where each ciphertext holds only one value and incurs significant overhead, HW layout is more efficient, packing $h\cdot w$ values per ciphertext, with $h$ and $w$ denoting input height and width. Figure 1 shows this for a $3\times 3\times 3$ input with two $3\times 2\times 2$ filters (stride 1, no padding), illustrating slot arrangements and gaps that may appear after convolution or other kernel operations. These gaps could be removed by remapping, but that requires additional multiplications and increases multiplicative depth. Instead, we use an adaptive (lazy) mapping strategy: some slots are left unused, and subsequent weights, biases, and coefficients are aligned with the slot arrangement of the preceding layer’s output. HW layout does not always fully utilize slots, especially when input channels have far fewer values than available slots. More compact layouts such as CHW layout (Dathathri et al., 2019) and its variants (Lee et al., 2022) address this by mapping multiple channels into one ciphertext, which reduces ciphertext counts and decreases the number of multiplications and additions, ultimately leading to faster inference. However, PPML relies on polynomial approximations that reduce accuracy compared to plaintext ReLU models. For this reason, we retain the less efficient HW layout, repurposing unused slots to improve accuracy instead of minimizing inference time. To offset the added latency, we introduce a parameter clustering technique.
@@ 4.2 Clustering of Convolution Parameters
Convolution layers account for most parameters in deep neural networks, and their weight handling under FHE is costly since each weight must be repeatedly encoded to match varying ciphertext layouts and levels. As these depend on kernel position and layer structure, the same scalar weight is often redundantly encoded. Pre-encoding all weight–layout combinations would lead to prohibitive memory usage, while on-demand encoding reduces memory but adds heavy computational overhead. To mitigate this, weights are limited to a small fixed set of representative values. Each representative is encoded once into a plaintext codebook, and during inference, the appropriate plaintext is retrieved rather than recomputed. Time and memory then scale with codebook size instead of the number of weights. Clustering provides a way to build this representative set by grouping similar weights and replacing each with its closest centroid, bounding the number of unique encodings.
@@
Formally, let a polynomial network contain $L$ convolution layers. The weight tensor of layer $l$ is $W^{(l)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$, where $O_{l}$ is the number of output channels, $I_{l}$ the number of input channels, and $H_{l}\times W_{l}$ the kernel size. Flattening all convolution weights gives $\theta=(\theta_{1},\dots,\theta_{P})\in\mathbb{R}^{P},$ where $P$ is the total number of parameters. To compress the model, each $\theta_{j}$ is replaced by its nearest codebook value from $\mathcal{C}=\{c_{1},\dots,c_{k}\}\subset\mathbb{R}$, where $k\ll P$ is the number of representatives in the codebook. The codebook entries act as centroids approximating the original weights. The choice of $k$ determines the efficiency-accuracy tradeoff: smaller $k$ reduces the number of encodings but increases approximation error, while larger $k$ produces finer approximation at higher computational and memory cost. Given a distance function $d:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ (assumed to be non-negative), each parameter is quantized as: $\tilde{\theta}_{j}=\arg\min_{c\in\mathcal{C}}d(\theta_{j},c)\text{ for all }j=1,\dots,P,$ producing a quantized parameter vector $\tilde{\theta}\in\mathcal{C}^{P}$. Quantization is applied elementwise, so each convolutional kernel retains its shape while its values are drawn from the codebook. The choice of clustering strategy determines how the codebook is constructed and how assignments are made.
@@
Full Clustering. A straightforward strategy is to cluster the entire parameter vector $\theta\in\mathbb{R}^{P}$ using a single global codebook $\mathcal{C}$. All convolution parameters, regardless of layer or kernel position, are quantized to the same shared set of representative values. This produces a uniform quantization scheme with easy implementation and a small codebook size.
@@
Limitation. While full clustering reduces the number of distinct weight values, it does not eliminate repeated plaintext encodings. Each weight must still match the ciphertext layout and level of its layer, so the same scalar value appearing in different layers must be re-encoded. Even within a single convolution, all weights share the same level, but weights from different filter columns map to different layouts and require separate encodings. Only weights in the same column across layer filters align in layout and level. We define this group as a slice, which can share plaintext encoding. Thus, full clustering reduces codebook size but not redundancy from mismatched levels and layouts.
@@
Slice Clustering. To address the redundancy remaining under full clustering, we refine quantization to the slice level. Since only weights within the same slice can share plaintext encodings, clustering is applied per slice rather than across all parameters. For a given layer $l$ and spatial position $s\in\{1,\dots,W_{l}\}$, the slice is $S^{(l)}_{s}=\{W^{(l)}_{o,i,h,s}|o=1,\dots,O_{l};\ i=1,\dots,I_{l};\ h=1,\dots,H_{l}\},$ which can be viewed as a vector in $\mathbb{R}^{O_{l}I_{l}H_{l}}$ (see §F.1 for an illustration). Each slice $S^{(l)}_{s}$ is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}\subset\mathbb{R}$, and quantization is applied elementwise: $\tilde{W}^{(l)}_{o,i,h,s}=\arg\min_{c\in\mathcal{C}^{(l)}_{s}}d(W^{(l)}_{o,i,h,s},c).$ This ensures clustering respects FHE structure while adapting to the local weight distribution of each slice. The additional storage from slice-specific codebooks is modest compared to the resulting gains in accuracy and efficiency.
@@ 4.3 Ensemble of Polynomial Network
Inference with a single polynomial network $g(x)$ often shows high variance across training runs. Let $g_{(m)}(x)\in\mathbb{R}^{K}$ be the logits from the $m$-th independently trained instance. For a fixed input, these logits can differ significantly, especially near decision boundaries, due to (1) approximation error from the fixed-point polynomial activation $p_{d}(\cdot)$ and (2) training stochasticity such as random weight initialization and mini-batch ordering. As a result, different $g_{(m)}$ may predict different classes for the same input. To reduce this variance, we use an ensemble of $M$ polynomial networks $\{g_{(m)}\}_{m=1}^{M}$. Each has the same architecture and activation $p_{d}(\cdot)$ but is trained independently with different seeds and mini-batch orders. For input $x$, the ensemble output is the average of logits: $\overline{g}(x)=\frac{1}{M}\sum_{m=1}^{M}g_{(m)}(x)$ with the predicted class $\hat{y}=\arg\max_{k\in\{1,\dots,K\}}\overline{g}(x)_{k}.$ Averaging smooths training noise and polynomial approximation errors, producing more stable and accurate predictions. Each $g_{(m)}$ is trained with the regularized loss $\mathcal{L}_{\mathcal{B}}$ (Eq. 1), ensuring compatibility with fixed-point polynomial inference under FHE. Crucially, this ensemble design adds no extra computation or memory overhead. As noted in §4.1, our HW layout leaves some ciphertext slots unused. We fill these slots with weights from different models, while all ensemble members share the same ciphertexts for inputs and activations. This reuse of ciphertext–plaintext structures avoids redundant ciphertexts and repeated encodings, making ensemble inference practical in the FHE framework.
@@
Limits of Ensemble with Clustering. While ensemble inference reuses unused ciphertext slots, its combination with parameter clustering creates a challenge. In slice clustering, each slice $S_{s}^{(l)}$ with $O_{l}\times I_{l}\times H_{l}$ weights is represented by $k$ centroids in codebook $\mathcal{C}_{s}^{(l)}$, reducing plaintext encodings to $k$ per slice. In an ensemble, however, each plaintext must encode parameters from all $M$ models. As centroids are unlikely to align across models, the count of unique encodings per slice grows as $\mathcal{O}(k^{M})$, canceling the gains of clustering and making the naive combination impractical.
@@
Slice Ensemble Clustering. To address the inefficiency of independent clustering, we extend slice clustering to ensembles by enforcing shared representatives across models. We cluster weights jointly at the same kernel position so that all models use a common codebook. Formally, the ensemble has $M$ polynomial networks. For convolution layer $l$, model $m$ has weights $W^{(l,m)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$. For kernel position $s\in\{1,\dots,W_{l}\}$, we extract slices: $S^{(l,m)}_{s}=\{W^{(l,m)}_{o,i,h,s}|o=1,\dots,O_{l};i=1,\dots,I_{l};h=1,\dots,H_{l}\}.$ Stacking slices across $M$ models produces $X^{(l)}_{s}=[S^{(l,1)}_{s}\;S^{(l,2)}_{s}\cdots S^{(l,M)}_{s}]\in\mathbb{R}^{N_{s}\times M},$ where $N_{s}=O_{l}I_{l}H_{l}$. Each row is an $M$-dimensional vector for the same weight coordinate across models (see §F.2 for an illustration). We cluster rows in $\mathbb{R}^{M}$: $\min_{\mathcal{C}^{(l)}_{s}\subset\mathbb{R}^{M}}\sum_{j=1}^{N_{s}}\min_{c\in\mathcal{C}^{(l)}_{s}}d(X^{(l)}_{s,j},c),$ with codebook $\mathcal{C}^{(l)}_{s}=\{c_{1},\dots,c_{k}\}$. Each $X^{(l)}_{s,j}$ is replaced by its nearest centroid, producing quantized slices $\tilde{X}^{(l)}_{s,j}\in\mathcal{C}^{(l)}_{s}$. Quantized weights $\tilde{W}^{(l,m)}$ are reconstructed by reshaping slices. This shared clustering aligns weights across models, mapping small variations at the same coordinate to a common centroid. It thus requires fewer distinct encodings, reducing codebook size and inference time while retaining the accuracy gains of ensemble.

### v2

@@ 4.3 Tower Reuse
In CKKS, the RNS moduli are typically selected as primes close to the scaling factor $\Delta$. This choice is motivated by the fact that the scaling factors of operands multiply after each homomorphic multiplication, yielding a ciphertext with scale $\Delta^{2}$. To prevent exponential growth of the scale, the rescale operation is used to reduce the scale back to $\Delta$, consuming one modulus from the modulus chain in the process. For this reason, the number of moduli depends on the application’s multiplicative depth $\delta$. Since homomorphic operations are carried out across all moduli, their number directly impacts computational efficiency. Rescaling also introduces approximation errors: the further the dropped modulus from $\Delta$, the larger the error. This issue is especially relevant for small scaling factors, as fewer RNS primes lie relatively close to $\Delta$. These errors can accumulate and amplify through subsequent homomorphic operations, increasing the likelihood of incorrect results. Using a large $\Delta$ has drawbacks, as it requires larger RNS moduli and therefore increases the ciphertext modulus $Q=\prod_{i=0}^{\mathcal{L}}q_{i}$, where $\mathcal{L}$ is the number of levels and $q_{i}$ are the RNS primes, which reduces the security level. One can increase security by increasing the polynomial degree $N$. However, it is undesirable because operating on larger polynomials slows computation. Thus, for a fixed security level and polynomial degree, there is a maximum allowable $Q$. Consequently, for an application-defined $\delta$, this imposes a bound on the maximum $\Delta$. In practice, when $\delta$ is large, as is common in deep learning models, the resulting $\Delta$ is small.
@@
Proposed Solution. We introduce tower reuse, a more general method for determining the RNS moduli by introducing sublevels within each level. Instead of enforcing $q_{i}\approx\Delta$, we allow $q_{i}\approx\Delta^{\ell}$, where $\ell$ denotes the number of sublevels. Rescaling is performed only when a ciphertext scale exceeds the sublevel capacity of the modulus to be dropped. Formally, $\lambda(x)>\lambda(q_{i})\implies\downarrow x$, where $x$ represents a ciphertext, $\downarrow$ denotes the rescale operation, and $\lambda(x)=\left\lfloor\log_{\Delta}{\Delta_{x}}\right\rceil$ returns the sublevel of a ciphertext, plaintext, or modulus: with $\Delta_{x}$ referring to the scale of $x$ and $\Delta$ being the default scaling factor defined by the encryption parameters.
Example. Suppose following an activation function, there is a convolution. Let modulus $q_{i}\approx\Delta^{2}$, thus $\lambda(q_{i})=2$. Consider an input $x$ with scaling factor $\Delta_{x}=\Delta$, giving $\lambda(x)=1$. Applying a quadratic activation $y=x^{2}+c_{1}x+c_{0}$ yields $\lambda(y)=2$, with $\lambda(c_{1})=1$ and $\lambda(c_{0})=2$. Next, a convolution $z=\sum_{i}\omega_{i}y_{i}+\alpha$ is performed. Since $\lambda(y)=2$, by setting $\lambda(\omega)=1$, we obtain $\lambda(z)=3$. As $\lambda(z)>\lambda(q_{i})$, a rescale is triggered. The rescale operation drops $q_{i}$ from the modulus chain, reducing the level $\Lambda(\cdot)$ by one and resetting the sublevel to one:
Tower reuse is an automatic inference-time technique governed by $q_{i}$, $\Delta$, and the network structure. It improves computational efficiency and reduces approximation errors. The efficiency gains arise from decreasing the number of RNS moduli by a factor of $\ell$, since the method allows $\ell$ levels of multiplication to be performed before rescaling, and from the corresponding reduction in the number of required rescale operations. The reduction in approximation error is likewise due to the fewer rescale operations, and also to the fact that the method permits the use of smaller $\Delta$ values together with larger $q_{i}$, making the moduli relatively closer to $\Delta^{\ell}$.
@@ 4.4 Impact on RNS Levels
The combination of our polynomial activation function with the structural optimizations proposed in this section reduces the number of RNS levels $\mathcal{L}$ required for multiplication compared to PILLAR [11], which previously achieved the lowest $\mathcal{L}$: from 79 to 16 for VGG-16, 87 to 18 for ResNet-18, 97 to 20 for ResNet-20, and 157 to 32 for ResNet-32.
Level Analysis. Table 1 summarizes the number of levels $\mathcal{L}$ required for multiplication across VGG and ResNet models under various optimization techniques. To illustrate, we focus on ResNet-18, which comprises 17 convolutional layers, batch normalization layers, and activation functions, along with a single average pooling and linear layer:
  • $P_{4}$. Prior work with the lowest $\mathcal{L}$ [11] uses a degree-4 polynomial activation, giving $\mathcal{L}=87$: convolutions, batch normalizations, average pooling, and linear layers each contribute $\mathcal{L}=1$, and the activation contributes $\mathcal{L}=3$.
  • $P_{2}$. By replacing the activation function with a degree-2 polynomial, as described in §3, the required level for activations drops to 2, reducing the model $\mathcal{L}$ to 70.
  • $P_{2}F$. Employing node fusing (§4.1) on $P_{2}$ removes all batch normalizations, which lowers $\mathcal{L}$ to 53.
  • $P_{2}R$. Applying weight redistribution (§4.2) to $P_{2}$ instead yields $\mathcal{L}=35$. The highest-order coefficients in batch normalizations and polynomial activations are normalized to one, simplifying them to $x+b_{0}$ ($\mathcal{L}=0$) and $x^{2}+c_{1}x+c_{0}$ ($\mathcal{L}=1$); the average-pooling divisor is likewise normalized to one ($\mathcal{L}=0$). These simplifications give $\mathcal{L}=35$.
  • $P_{2}FR$. Combining node fusing and weight redistribution retains $\mathcal{L}=35$, but reduces homomorphic operations, since node fusing eliminates batch normalizations.
  • $P_{2}FRT$. Applying tower reuse (§4.3) to $P_{2}FR$ lets a nonlinear polynomial activation and subsequent linear kernel share a level, halving the number of levels to $\mathcal{L}=18$, following $\mathcal{L}_{P_{2}FRT}=\lceil\frac{\mathcal{L}_{P_{2}FR}}{2}\rceil$ for VGG and ResNet models.
In summary, by combining a lower-degree polynomial activation with structural optimizations, we reduce levels by nearly a factor of five compared to prior work.
Table 1: Number of RNS levels required for homomorphic multiplications in VGG and ResNet models. $P_{4}$: Degree-4 polynomial activation used in PILLAR [11]. $P_{2}$: Degree-2 polynomial activation (§3). $P_{2}F$: $P_{2}$ with node fusing (§4.1). $P_{2}R$: $P_{2}$ with weight redistribution (§4.2). $P_{2}FR$: $P_{2}$ with node fusing and weight redistribution. $P_{2}FRT$: $P_{2}FR$ with tower reuse (§4.3).
Model $P_{4}$ $P_{2}$ $P_{2}F$ $P_{2}R$ $P_{2}FR$ $P_{2}FRT$
VGG-16 79 64 51 31 31 16
ResNet-18 87 70 53 35 35 18
ResNet-20 97 78 59 39 39 20
ResNet-32 157 126 95 63 63 32
5 Co-design Techniques
@@ 5.1 Data Layout
Data layout defines how model inputs and weights are mapped into FHE ciphertext and plaintext slots during encoding. In this work, we adopt the HW layout [9], which assigns each input channel to a separate ciphertext and fills its slots with spatial values indexed by two–dimensional coordinates. Unlike the naive layout [13], where each ciphertext holds only one value and incurs significant overhead, HW layout is more efficient, packing $h\cdot w$ values per ciphertext, with $h$ and $w$ denoting input height and width. Figure 4 shows this for a $3\times 4\times 4$ input with a $3\times 2\times 2$ filters (stride 1, no padding), illustrating slot arrangements and gaps that may appear after convolution or other kernel operations. These gaps could be removed by remapping, but that requires additional multiplications and increases multiplicative depth. Instead, we use an adaptive (lazy) mapping strategy: some slots are left unused, and subsequent weights, biases, and coefficients are aligned with the slot arrangement of the preceding layer’s output. HW layout does not always fully utilize slots, especially when input channels have far fewer values than available slots. More compact layouts such as CHW layout [9] and its variants [20] address this by mapping multiple channels into one ciphertext, which reduces ciphertext counts and decreases the number of multiplications and additions, ultimately leading to faster inference. However, PPML relies on polynomial approximations that reduce accuracy compared to plaintext ReLU models. For this reason, we retain the less efficient HW layout, repurposing unused slots to improve accuracy instead of minimizing inference time. To offset the added latency, we introduce a parameter clustering technique.
Figure 4: Illustration of HW layout for convolution on a $3\times 4\times 4$ input and a $3\times 2\times 2$ filter, configured with padding 0 and stride 1.
@@ 5.2 Clustering of Convolution Parameters
Convolution layers account for the majority of parameters in deep neural networks, and their handling under FHE is particularly costly. Each convolution weight must be encoded repeatedly to match varying ciphertext layouts and levels, which depend on kernel position and layer structure. As a result, the same scalar weight is often redundantly encoded many times. Pre-encoding all possible weight–layout combinations would lead to prohibitive memory usage, while encoding weights on-demand reduces memory usage but introduces heavy computational cost. To mitigate this overhead, we restrict weights to a small, fixed set of representative values. Each representative is encoded once into a plaintext codebook, and during inference, the appropriate plaintext is retrieved rather than recomputed. Consequently, both runtime and memory scale with the codebook size rather than with the total number of weights. Clustering provides a mechanism for constructing such a representative set by grouping similar weights and replacing each with its nearest centroid, thereby bounding the number of unique encodings.
@@
Formally, consider a polynomial network with $L$ convolution layers. The weight tensor of layer $l$ is denoted by $W^{(l)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$, where $O_{l}$ is the number of output channels, $I_{l}$ the number of input channels, and $H_{l}\times W_{l}$ the kernel size. Flattening all convolution weights gives $\theta=(\theta_{1},\dots,\theta_{P})\in\mathbb{R}^{P},$ where $P$ is the total number of convolution parameters. To compress the model, each $\theta_{j}$ is replaced by its nearest codebook value from $\mathcal{C}=\{c_{1},\dots,c_{k}\}\subset\mathbb{R}$, where $k\ll P$ is the number of representatives in the codebook. The codebook entries act as centroids approximating the original weights. The choice of $k$ determines the efficiency-accuracy tradeoff: smaller $k$ reduces the number of encodings but increases approximation error, while larger $k$ produces finer approximation at higher computational and memory cost. Given a distance function $d:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ (assumed to be non-negative), each parameter is quantized as: $\tilde{\theta}_{j}=\arg\min_{c\in\mathcal{C}}d(\theta_{j},c)\text{ for all }j=1,\dots,P,$ producing a quantized parameter vector $\tilde{\theta}\in\mathcal{C}^{P}$. Quantization is applied elementwise, so each convolutional kernel retains its shape while its values are drawn from the codebook. The choice of clustering strategy determines how the codebook is constructed and how assignments are performed.
@@
Full Clustering. A straightforward strategy is to cluster the entire parameter vector $\theta\in\mathbb{R}^{P}$ using a single global codebook $\mathcal{C}$. All convolution parameters, regardless of layer or kernel position, are quantized to the same shared set of representative values. This produces a uniform quantization scheme with simple implementation and a small codebook size.
@@
Limitation. While full clustering reduces the number of distinct weight values, it does not eliminate redundant plaintext encodings. Each weight must still match the ciphertext layout and level of its layer, so the same scalar values appearing in different layers must be re-encoded. Even within a single convolution layer, although all weights share the same level, weights in different filter columns map to different layouts and thus require separate encodings. Only weights in the same column across layer filters align in layout and level. We refer to such a group as a slice, which can share plaintext encoding. Thus, full clustering reduces codebook size but fails to address redundancy arising from mismatched levels and layouts.
@@
Slice Clustering. To address the redundancy remaining under full clustering, we refine quantization at the slice level. Since only weights within the same slice can share plaintext encodings, clustering is applied independently to each slice rather than across all parameters. For a given layer $l$ and spatial position $s\in\{1,\dots,W_{l}\}$, we define the slice $S^{(l)}_{s}=\{W^{(l)}_{o,i,h,s}|o=1,\dots,O_{l};\ i=1,\dots,I_{l};\ h=1,\dots,H_{l}\},$ which can be viewed as a vector in $\mathbb{R}^{O_{l}I_{l}H_{l}}$. Each slice $S^{(l)}_{s}$ is clustered independently using its own codebook $\mathcal{C}^{(l)}_{s}\subset\mathbb{R}$, and quantization is applied elementwise: $\tilde{W}^{(l)}_{o,i,h,s}=\arg\min_{c\in\mathcal{C}^{(l)}_{s}}d(W^{(l)}_{o,i,h,s},c).$ This ensures clustering respects FHE structure while adapting to the local weight distribution of each slice. Figure 5 illustrates slice clustering for a single convolution layer. Convolution filters are decomposed into vertical slices along the kernel width dimension $W_{l}$. Each colored slice $S^{(l)}_{s}$ is clustered independently using its own codebook. Each three-dimensional block represents a convolution filter associated with one output channel. The three axes correspond to kernel height $H_{l}$ (vertical), kernel width $W_{l}$ (horizontal), and input channels $I_{l}$ (depth). The collection of such filters forms the output channels $O_{l}$. Slices are taken as vertical strips along the width dimension $W_{l}$, and the figure highlights them with distinct colors. The blue strip corresponds to slice $S^{(l)}_{1}$, the yellow strip to slice $S^{(l)}_{2}$, and the pink strip to slice $S^{(l)}_{3}$. Each slice $S^{(l)}_{s}$ groups together all weights $W^{(l)}_{o,i,h,s}$} at a fixed width position $s$ across output channels $O_{l}$, input channels $I_{l}$, and kernel height $H_{l}$, and is clustered independently with its own codebook $\mathcal{C}^{(l)}_{s}$. This illustration emphasizes that clustering is performed slice by slice, rather than across the entire filter volume.
Figure 5: Illustration of single model slice clustering. Convolution filters are decomposed into slices along the kernel width, and each slice is clustered independently with its own codebook.
@@ 5.3 Ensemble of Polynomial Network
Inference with a single polynomial network $g(x)$ often shows high variance across training runs. Let $g_{(m)}(x)\in\mathbb{R}^{K}$ denote the logits by the $m$-th independently trained instance. For a fixed input, these logits can differ significantly, especially near decision boundaries, due to (1) approximation error introduced by the fixed-point polynomial activation $p_{d}(\cdot)$ and (2) training stochasticity such as random weight initialization and mini-batch ordering. As a result, different $g_{(m)}$ instances may predict different classes for the same input. To reduce this variance, we use an ensemble of $M$ polynomial networks $\{g_{(m)}\}_{m=1}^{M}$. All ensemble members share the same architecture and activation $p_{d}(\cdot)$, but are trained independently with different random seeds and mini-batch orders. Given an input $x$, the ensemble output is defined as the average of logits: $\overline{g}(x)=\frac{1}{M}\sum_{m=1}^{M}g_{(m)}(x)$ with the predicted class $\hat{y}=\arg\max_{k\in\{1,\dots,K\}}\overline{g}(x)_{k}.$ Logit averaging smooths both training noise and polynomial approximation errors, producing more stable and accurate predictions. Each $g_{(m)}$ is trained using the same regularized loss $\mathcal{L}_{\mathcal{B}}$ (Eq. 1), ensuring compatibility with fixed-point polynomial inference under FHE. Crucially, this ensemble design adds no additional computation or memory overhead. As described in §5.1, our HW layout leaves a subset of ciphertext slots unused. We populate these slots with weights from different ensemble members, while all models share the same ciphertexts for inputs and activations. This reuse of ciphertext–plaintext structures avoids redundant ciphertexts and repeated encodings, making ensemble inference practical within the FHE framework.
@@
Limits of Ensemble with Clustering. While ensemble inference reuses unused ciphertext slots, naively combining it with parameter clustering introduces a challenge. In slice clustering, each slice $S_{s}^{(l)}$ containing $O_{l}\times I_{l}\times H_{l}$ weights is represented by $k$ centroids in codebook $\mathcal{C}_{s}^{(l)}$, reducing the number of plaintext encodings to $k$ per slice. In an ensemble setting, however, each plaintext must encode parameters from all $M$ models. Since centroids are unlikely to align across independently trained models, the number of distinct encodings per slice grows as $\mathcal{O}(k^{M})$, negating the benefits of clustering and making the naive combination impractical.
@@
Slice Ensemble Clustering. To address the inefficiency of independent clustering, we extend slice clustering to ensemble models by enforcing a shared set of representatives across all ensemble members. Specifically, we cluster weights jointly at the same kernel position so that all models use a common codebook. Consider an ensemble of $M$ polynomial networks. For convolutional layer $l$, model $m$ has weights $W^{(l,m)}\in\mathbb{R}^{O_{l}\times I_{l}\times H_{l}\times W_{l}}$. For each kernel width position $s\in\{1,\dots,W_{l}\}$, we extract slices: $S^{(l,m)}_{s}=\{W^{(l,m)}_{o,i,h,s}|o=1,\dots,O_{l};i=1,\dots,I_{l};h=1,\dots,H_{l}\}.$ Stacking slices from all $M$ models produces $X^{(l)}_{s}=[S^{(l,1)}_{s}\;S^{(l,2)}_{s}\cdots S^{(l,M)}_{s}]\in\mathbb{R}^{N_{s}\times M},$ where $N_{s}=O_{l}I_{l}H_{l}$. Each row of $X^{(l)}_{s}$ is an $M$-dimensional vector corresponding to the same weight coordinate across all models. We cluster these rows in $\mathbb{R}^{M}$ by solving: $\min_{\mathcal{C}^{(l)}_{s}\subset\mathbb{R}^{M}}\sum_{j=1}^{N_{s}}\min_{c\in\mathcal{C}^{(l)}_{s}}d(X^{(l)}_{s,j},c),$ where $\mathcal{C}^{(l)}_{s}=\{c_{1},\dots,c_{k}\}$ is the shared codebook. Each row $X^{(l)}_{s,j}$ is replaced by its nearest centroid, producing quantized slices $\tilde{X}^{(l)}_{s,j}\in\mathcal{C}^{(l)}_{s}$. Quantized weights $\tilde{W}^{(l,m)}$ are then reconstructed by reshaping the corresponding slices. This shared clustering aligns weights across ensemble members, mapping small per-model variations at the same coordinate to a common centroid. It thus requires fewer distinct encodings, reducing codebook size and inference time while retaining the accuracy gains of ensembling. Figure 6 illustrates how slice clustering is extended from a single model to an ensemble of polynomial networks.
Figure 6: Illustration of slice-wise clustering for an ensemble. Slices at the same kernel width position from different models are stacked and clustered jointly using a shared codebook.
Each three-dimensional block in the figure represents the convolution filters of a single model instance, spanning kernel height $H_{l}$, kernel width $W_{l}$, and input channels $I_{l}$. The collection of such filters across output channels $O_{l}$ forms the layer weights $W^{(l,m)}$ for model $m$. For a fixed layer $l$, slices are extracted as vertical strips along the width dimension $W_{l}$, with colors indicating different slices $S^{(l,m)}_{s}$ within each model. Unlike the single-model case, however, we now align slices from multiple models at the same width position $s$ and stack them together. This produces a slice matrix $X^{(l)}_{s}\in\mathbb{R}^{N_{s}\times M}$, where each row corresponds to one weight coordinate across all $M$ models, and $N_{s}=O_{l}I_{l}H_{l}$. Clustering is then performed jointly in $\mathbb{R}^{M}$ across models, producing a shared codebook $\mathcal{C}^{(l)}_{s}$ for slice $s$. The illustration emphasizes that ensemble clustering is not applied to each model independently. Instead, slices are stacked across models at the same spatial position, and clustering is performed jointly.